k islargest positive integer such that 2009^k divides factorial 2010 exactly. find k

Grade:12

3 Answers

jitender lakhanpal
12 years ago

Dear manoj,

 there cant be any value of k for which 2009k can be divided by 2010

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender 

manoj jain
12 years ago

solution is not according to question

manoj jain
12 years ago

2009 =7.7.41

2010! = 1.2.3............2010

        =   ( 7.14............82terms)(41.82............49terms) .others factors

        =    7^82(1.2.3............82)41^49(1.2.......49).other factors

        = 7^82  ( 7.14......11terms) (other factors )41^49(41)(other factors)  

         = 7^82.7^11.41^50.other factors     

         =( 7.7)^50.41 ^50 .otherfactors

        =(7.7.41)^50.other factors

         =(2009)50.other factrs

 k=50 ans